OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x.

OpenStudy (anonymous):

f(x) = 2/x and g(x) = 2/x

OpenStudy (amorfide):

f(g(x)) well if you had f(3) you would substitute 3 into where you find x so you have f(g(x)) so substitute g(x) where you find x in f(x) this means you are substituting the function g(x) into f(x)

Parth (parthkohli):

\[\rm f(g(x)) = {2 \over \large {2 \over x}}\]and\[\rm g(f(x)) = {2 \over \large{2 \over x}}\]Simplify.

OpenStudy (amorfide):

@ParthKohli PARTH YOU ARE ALIVE

Parth (parthkohli):

@amorfide YA I AM BACK FROM THE DEAD

OpenStudy (anonymous):

ummm x for both???

OpenStudy (amorfide):

showing that f(g(x)) = x and g(f(x)) = x. that is exactly what you should get well done

OpenStudy (anonymous):

thanks! (that was easier than I thought) ^^